Circular No. 7495

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dm IN UGC 11198 A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery of an apparent supernova (mag ∼ 16.1) on an unfiltered image taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Sept. 24.1 UT. SN 2000dm, which was confirmed on an earlier image taken on Sept. 16.1 at mag ∼ 18.5, is located at α = 18^{h}18^{m}58^{s}.97, δ = +16^{o}14′53′′.4 (equinox 2000.0), which is 3′′.6 west and 4′′.7 south of the nucleus of UGC 11198. A KAIT image of the same field taken on Sept. 11.1 showed nothing at the position of the new object (limiting mag ∼ 19.0).

V1333 AQUILAE R. Jain, C. Bailyn, and P. Coppi, Yale University; M. Garcia, Harvard-

Smithsonian Center for Astrophysics; A. Levine, Massachusetts Institute of Technology; J. Orosz, Utrecht University; P. Lu, Western Connecticut State University; and J. Espinoza and D. Gonzalez, Cerro Tololo Interamerican Observatory (CTIO), report on YALO consortium observations: “Aquila X-1 = V1333 Aql appears to be beginning a new optical outburst. This x-ray binary outbursts approximately once per year (e.g., 1999 May 8.3 UT; cf. IAUC 7161), and was due to erupt. We have been monitoring this source since 2000 Apr. 11 in R using the YALO 1-m telescope at CTIO. Compared to the quiescent average magnitude, which has a fluctuation < 0.1 mag, the source had brightened in R by 0.45 ± 0.02 mag on Sept. 22.15. The R-band brightness continued to increase, reaching 0.62 ± 0.02 and 0.71 ± 0.03 mag above quiescence on Sept. 23.01 and 24.17, respectively. The amplitude of this outburst is larger than the peak from recent mini-outbursts (IAUC 7429, 7423, 7017). We will continue monitoring this source at optical, infrared, and x-ray (using RXTE) wavelengths. Observations at all wavelengths are encouraged.”

SUPERNOVA 2000db IN NGC 3949 CCD R magnitudes from K. Hornoch, Lelekovice, Czech Republic:

Aug. 20.823 UT, 13.7; 26.804, 13.8; 27.796, 13.7; 29.810, 13.9; Sept. 9.788, 14.2; 10.791, 14.3; 23.758, 14.4. CCD V magnitude from D. Hanžl, Brno, Czech Republic: Aug. 29.82, 14.7 ± 0.1.

2000 September 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green